OpenStudy (anonymous):

use rational exponents to write single radical expression: ^4√8 * ^2√7

jimthompson5910 (jim_thompson5910):

First, use the fact that \[\sqrt[n]{x^m}=x^{\frac{m}{n}}\] to rewrite \[\sqrt[4]{8}\] into \[8^{\frac{1}{4}}\] and to rewrite \[\sqrt[2]{7}\] into \[7^{\frac{1}{2}}\] So after all that, you now have the expression \[8^{\frac{1}{4}}7^{\frac{1}{2}}\] see what you can do with that.

OpenStudy (anonymous):

jim_thompson5910 i'm not understanding

jimthompson5910 (jim_thompson5910):

I'm converting radical expressions (expressions with roots) into expressions that involve exponents only.